I have been working on an assignment that involves solving using Cramer's rule I have set up the matrices but cannot get any further.
2x- y+ z =0 D= 2 -1 1 Dx 0 -1 +1
4x+6y-2z =0 4 +6 -2 0 +6 -2
x-2y- z =-9 1 -2 -1 -9 -2 -1
Dy 2 0 1 Dz 2 -1 0
4 0 -2 4 6 0
1 -9 -1 1 -2 -9
What the heck do I do now?

Answer by stanbon(48568)

Calculate the determinant of D=-36, Dx=36, Dy=-72, Dz-144
-----------
Then Cramer determined the following:
x = Dx/D , by this I mean the quotient of their determinants
y = Dy/D , same
z = Dz/D , same
-------------------
Solution:
x =-1
y = 2
z = 4
===============
Cheers,
Stan H.